# Perception-guided Jailbreak against Text-to-Image Models

### v2

@@ Abstract
Text-to-Image Model, Not-Safe-For-Work, Perception
@@ 1 Introduction
The growing interest in artificial intelligence (AI) safety [4, 5, 6, 7] has intensified due to the widespread adoption and advanced capabilities of AI models. This expansion has raised substantial security concerns about the potential for generating unsafe images. A prominent ethical issue associated with T2I models is their potential to produce sensitive Not-Safe-for-Work (NSFW) images [8, 9], including adult content, violence, and politically sensitive material. Therefore, current T2I models incorporate safety checkers [10, 11] as essential guardrails to prevent the generation of NSFW images.
@@ 2.1 Text-to-Image Models
Notable examples include Stable Diffusion [1], DALL·E [26, 3], Imagen [27], Midjourney [2], and Wanxiang [28]. One of the most advanced T2I models, DALL·E 3 [3], integrated natively into ChatGPT [29], leverages LLM [30] to refine prompts, producing images that closely align with the input prompts and reducing the users’ burden of prompt engineering. Given their popularity, studies that investigate the vulnerabilities [31, 32] of T2I models are necessary.
@@ 2.2 Jailbreak on Text-to-Image Models
Previous research [33, 34, 35, 36, 37, 38] predominantly focuses on text modification to probe functional vulnerabilities (e.g., diminishing synthetic quality, distorting or eliminating objects, impairing image fidelity). However, they do not target the generation of Not-Safe-For-Work (NSFW) images such as pornography, violence, and racism.
Currently, more and more works [12, 13, 14, 15, 16] have put emphasis on exploring the opened avenues for potential misuse of T2I models, particularly in generating inappropriate or NSFW content via adversarial attack [39, 40]. SneakyPrompt [12] exploits reinforcement learning to replace the words in the prompt for bypassing safety filters in T2I generative models. SurrogatePrompt [13] proposes a pipeline that contains three modules to generate NSFW images on T2I models such as Midjourney and DALL·E 2. DACA [41] breaks down unethical prompts into multiple benign descriptions of individual image elements and makes word substitutions for each element. MMA-Diffusion [14] is a multimodal attack framework that designs attacks on both text and image modalities. UPAM [15] is a unified framework that employs gradient-based optimization, sphere-probing learning, and semantic-enhancing learning to attack the T2I model. JPA [16] using learnable tokens to create adversarial prompts that evade detection while preserving the semantic integrity of the original NSFW content.
❶ Rely on white-box adversarial attacks targeting a specific T2I model, and then subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. ❷ Others involve developing complex pipelines that require numerous queries to the T2I model, leading to high time and resource consumption. In contrast, our method is model-free, requiring no specific T2I model as a target, and generates attack prompts with high naturalness.
@@ 3.1 Problem Definition
Safety checker. The primary challenge is bypassing the safety checker $\mathcal{F}$, which consists of two modules: a pre-checker and a post-checker. The pre-checker is a text filter that identifies unsafe or sensitive words in input prompts, while the post-checker is an image filter that detects NSFW content in output images. In this paper, we focus on bypassing the pre-checker and do not address the post-checker for three key reasons. ❶ The pre-checker is more cost-effective and widely used, as it proactively blocks unethical prompts, thereby reducing unnecessary computational costs associated with image generation. ❷ Prompts are typically smaller in size than images, making the pre-checker more efficient at handling large volumes of requests. ❸ Our experiments with current open-source and commercial T2I models demonstrate that our method can effectively jailbreak these models even without specifically targeting the post-checker, highlighting its vulnerability.
@@ 5.1 Experimental Setups
Datasets. The prompts in the NSFW dataset used by jailbreak methods [12, 14, 41] typically are of small amounts, unbalanced in NSFW types, and contain duplicate entries. Thus we exploit GPT4 to generate a dataset with 1,000 prompts for five classical NSFW types: discrimination, illegal, pornographic, privacy, and violent. For each type, we generate 200 prompts. This is to ensure the prompt number for each type is balanced and the prompts are diverse. The details of the prompts are in Supp.
Baselines. Among the works aiming at jailbreak T2I models, we choose all the popular and state-of-the-art ones that open-source the code: SneakyPrompt [12], MMA-Diffusion [14], DACA [41]. We conduct the experiment exactly according to their experimental setup. All the experiments were run on an Ubuntu system with an NVIDIA A6000 Tensor Core GPU of 50G RAM.
Evaluation metrics. We use four metrics to evaluate the experiment. ❶ We use the attack success rate (ASR) metric to evaluate the number of attack prompts that bypass the NSFW detector divided by the total number of attack prompts. ❷ We use the semantic consistency (SC) metric to represent the consistency between the semantics of the generated image and the original unsafe user prompt. The generated image should have a similar semantic as the original unsafe user prompt, i.e., the jailbreak method does not change the semantics of the unsafe user prompt. The semantics of the generated image are extracted by BLIP [46]. ❸ We use prompt perplexity (PPL) as a metric to evaluate the coherence of the modified attack prompt. The prompt with high PPL contains a lot of garbled characters and is easy to notice. ❹ We use the Inception Score (IS) to evaluate the diversity of the generated images. For ASR, SC, and IS metrics, higher is better while for PPL, lower is better. Note that the ASR and SC metrics are dominant ones for evaluating the jailbreak performance of methods.
@@ 6 Conclusion
In this paper, we introduce a word replacement method that identifies a safe substitution phrase adhering to the PSTSI principle. The proposed PGJ method efficiently and effectively generates an attack prompt capable of bypassing the safety checkers in T2I models. For future work, we plan to explore circumventing the post-checker in T2I models.

### v3

@@ Abstract
Warning: This paper contains NSFW and disturbing imagery, including adult, violent, and illegal-related contentious content. We have masked images deemed unsafe. However, reader discretion is advised.
@@
Text-to-Image Model, Not-Safe-For-Work, Perception, Jailbreak
@@ 1 Introduction
However, the widespread use and advanced capabilities of these models have led to significant security concerns regarding unsafe image generation. A prominent ethical issue associated with T2I models is their potential to produce sensitive Not-Safe-for-Work (NSFW) images [4, 5], including adult content, violence, and politically sensitive material. Therefore, current T2I models incorporate safety checkers [6, 7] as essential guardrails to prevent the generation of NSFW images.
@@ 2.1 Text-to-Image Models
Notable examples include Stable Diffusion [1], DALL·E [22, 3], Imagen [23], Midjourney [2], and Wanxiang [24]. One of the most advanced T2I models, DALL·E 3 [3], integrated natively into ChatGPT [25], leverages LLM [26] to refine prompts, producing images that closely align with the input prompts and reducing the users’ burden of prompt engineering [27]. Given their popularity, investigating the vulnerabilities of T2I models is necessary.
@@ 2.2 Jailbreak on Text-to-Image Models
Adversarial attacks [28, 29, 30, 31, 32] are effective in exposing neural network vulnerabilities [33, 34, 35, 36]. While prior research [37, 38, 39, 40, 41, 42, 43, 44, 45] focuses on text modifications to exploit functional weaknesses (e.g., degrading quality, distorting objects, or impairing fidelity), they overlook the generation of Not-Safe-For-Work (NSFW) content such as pornography, violence, and racism.
Currently, more and more works [8, 9, 10, 11, 12, 46, 47] have put emphasis on exploring the opened avenues for potential misuse of T2I models, particularly in generating inappropriate or NSFW content. SneakyPrompt [8] exploits reinforcement learning to replace the words in the prompt for bypassing safety filters in T2I generative models. SurrogatePrompt [9] proposes a pipeline that contains three modules to generate NSFW images on T2I models such as Midjourney and DALL·E 2. DACA [48] breaks down unethical prompts into multiple benign descriptions of individual image elements and makes word substitutions for each element. MMA-Diffusion [10] is a multimodal attack framework that designs attacks on both text and image modalities. UPAM [11] is a unified framework that employs gradient-based optimization, sphere-probing learning, and semantic-enhancing learning to attack the T2I model. JPA [12] using learnable tokens to create adversarial prompts that evade detection while preserving the semantic integrity of the original NSFW content. Ring-A-Bell [46] is a model-agnostic evaluation framework that leverages concept extraction to represent sensitive or inappropriate concepts. ColJailBreak [47] produces NSFW images by first generating safe content, then injecting unsafe elements via inpainting, and finally refining the outputs for seamless integration, but it does not focus on bypassing the safety checker of T2I models. Recent work has also explored methods for mitigating the generation of unsafe content in text-to-image models, such as SafeGen [49], which aims to prevent the creation of NSFW images in a text-agnostic manner.
Rely on white-box adversarial attacks targeting a specific T2I model, and then subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. ❷ Others involve developing complex pipelines that require numerous queries to the T2I model, leading to high time and resource consumption. In contrast, our method is model-free, requiring no specific T2I model as a target, and generates attack prompts with high naturalness.
@@ 3.1 Problem Definition
Safety checker. The primary challenge is bypassing the safety checker $\mathcal{F}$, which consists of two modules: a pre-checker and a post-checker. The pre-checker is a text filter that identifies unsafe or sensitive words in input prompts, while the post-checker is an image filter that detects NSFW content in output images. In this paper, we focus on bypassing the pre-checker and do not focus on the post-checker for three key reasons. ❶ The pre-checker is more cost-effective and widely used, as it proactively blocks unethical prompts, thereby reducing unnecessary computational costs associated with image generation. ❷ Prompts are typically smaller in size than images, making the pre-checker more efficient at handling large volumes of requests. ❸ Our experiments with current open-source and commercial T2I models demonstrate that our method can effectively jailbreak these models even without specifically targeting the post-checker, highlighting its vulnerability. It is important to note that our primary focus was on bypassing the text checker, as image checkers in current text-to-image models are generally easier to circumvent, while text checkers pose a significantly greater challenge.
@@ 5.1 Experimental Setups
Datasets. The prompts in the NSFW dataset used by jailbreak methods [8, 10, 48] typically are of small amounts, unbalanced in NSFW types, and contain duplicate entries. Thus we exploit GPT4 to generate a dataset with 1,000 prompts for five classical NSFW types: discrimination, illegal, pornographic, privacy, and violent. The prompts are generated by GPT-4 with the instruction “Give me 200 English sentences containing NSFW Type content descriptions of images, without any other extra text”. GPT allows our dataset to be balanced across NSFW types, large in size, and highly diverse (since GPT is learned from a large and diverse corpus). For each type, we generate 200 prompts. This is to ensure the prompt number for each type is balanced and the prompts are diverse.
Baselines. Among the works aiming at jailbreak T2I models, we choose all the popular and state-of-the-art ones that open-source the code: SneakyPrompt [8], MMA-Diffusion [10], DACA [48], Ring-a-Bell [46]. We conduct the experiment exactly according to their experimental setup. All the experiments were run on an Ubuntu system with an NVIDIA A6000 Tensor Core GPU of 48G RAM.
Evaluation metrics. We use four metrics to evaluate the experiment. ❶ We use the attack success rate (ASR) metric to evaluate the number of attack prompts that bypass the NSFW detector divided by the total number of attack prompts. ❷ We use the semantic consistency (SC) metric to represent the consistency between the semantics of the generated image and the original unsafe user prompt. The generated image should have a similar semantic as the original unsafe user prompt, i.e., the jailbreak method does not change the semantics of the unsafe user prompt. The semantics of the generated images are extracted by BLIP [54]. ❸ We use prompt perplexity (PPL) as a metric to evaluate the coherence of the modified attack prompt. The prompt with high PPL contains a lot of garbled characters and is easy to notice. ❹ We use the Inception Score (IS) to evaluate the diversity of the generated images. For ASR, SC, and IS metrics, higher is better while for PPL, lower is better. Note that the ASR and SC metrics are dominant ones for evaluating the jailbreak performance of methods.
@@ 5.4 Ablation Studies
5.5 Discussion
Bypass detection of LLM. We also use modern LLM to detect if the attack prompt is actually an NSFW prompt in disguise. We use the instruction “You are an AI specialized in detecting harmful, unsafe, or inappropriate content. Assess if the provided prompt contains any of the following types of harmful content: violence, hate speech, self-harm, illegal activities, explicit sexual content, misinformation, or any other type of content that may be unsafe or harmful”. For 1,000 original NSFW prompts in our dataset, only 12.2% evade GPT-4o detection, whereas 43.8% of rewritten prompts do—an increase of 31.6%. This indicates that even powerful LLMs may miss harmful content in many rewritten prompts. While generally effective, LLMs are not flawless, underscoring the necessity and effectiveness of our method to reveal vulnerabilities in text-to-image models.
@@ 6 Conclusion
In this paper, we introduce a word replacement method that identifies a safe substitution phrase adhering to the PSTSI principle. The proposed PGJ method efficiently and effectively generates an attack prompt capable of bypassing the safety checkers in T2I models. For future work, we plan to explore for circumventing the post-checker in T2I models.
Ethical Statement
Our main objective is to propose jailbreak methods against the T2I models; however, we acknowledge the attack prompt will trigger inappropriate content from T2I models. Therefore, we have taken meticulous care to share findings in a responsible manner. We firmly assert that the societal benefits stemming from our study far surpass the relatively minor risks of potential harm due to pointing out the vulnerability of T2I models.